OpenStudy (anonymous):

Help me please! d(x)=x + pi Standard form: Polynomial Name: Degree: Leading Coefficient: Constant term:

OpenStudy (anonymous):

@Data_LG2 maybe you can help more here :)

OpenStudy (anonymous):

Standard form: y=mx+c Polynomial Name: Degree:1 Leading Coefficient: x Constant term:pi

OpenStudy (anonymous):

hmm.. standard forfm is Ax+By+C, arrange your equation in terms of this. d(x) is your y.. C=pi

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

@INeedHelpPlease? isn't that a slope-intercept form?

OpenStudy (anonymous):

d(x)=x + pi y = mx +c

OpenStudy (anonymous):

So wouldn't standard form just be x+pi

OpenStudy (anonymous):

\(y=x+ pi \) --> \(x-y +pi=0\)

OpenStudy (anonymous):

wait.. I'm confused... what exactly would be the standard form

OpenStudy (anonymous):

\(Ax + By +C = 0 \ is\ the\ standard\ form\ of\ the\ equation\) \(y=mx+b\ is\ the\ slope-intercept\ form\ of\ linear\ equation\)

OpenStudy (anonymous):

I know what standard form is, but I have to rearrange \[x+\pi \]

OpenStudy (anonymous):

yes, so what's the problem?

OpenStudy (anonymous):

well the answer is either x+pi or pi+x

OpenStudy (anonymous):

x−y+pi=0 --> this is the standard form already ?

OpenStudy (anonymous):

That's what I thought! So the equation is already in it's standard form of x+pi. What about the name?

OpenStudy (anonymous):

^that's the only one i don't know XD is it "LINEAR FUNCTION" , like that?

OpenStudy (anonymous):

I think they mean like quadratic binomial or such

OpenStudy (anonymous):

So linear binomial?

OpenStudy (anonymous):

linear is different from binomial ..

OpenStudy (anonymous):

linear for one exponent and binomial for two terms

OpenStudy (anonymous):

oh sweet! yeah, lol sorry i'm messing up today, i don't know what's happening to me.. so yeah, i think you're right(^_^)

OpenStudy (anonymous):

haha it's fine. What about degree, leading, and constant. I think leading is x and constant is pi right?

OpenStudy (anonymous):

leading coefficient is 1 because it's only "x", if it's "2x" leading coefficient will be 2

OpenStudy (anonymous):

Oh gotcha. What about constant and degree?

OpenStudy (anonymous):

IneedHElp, already did that part XD , scroll up, i think he/she is right (^_^)

OpenStudy (anonymous):

IneedHelp said leading is x

OpenStudy (anonymous):

i mean he is right for the constant and degree

OpenStudy (anonymous):

Oh ok haha

OpenStudy (anonymous):

anything else? (^_^)

OpenStudy (anonymous):

Yeah haha this is a pretty long review. Should I start a new question or post here?

OpenStudy (anonymous):

lol, XD